# Radio Observations of a Candidate Redback Millisecond Pulsar: 1FGL J0523.5-2529

### v1

@@ 5 Conclusions
Acknowledgements.
The authors acknowledge the insights and comments from Jules Halpern, Karen Perez, Wael Farah, Colin Clark, Scott Ransom, Lawrence Toomey, Joe Glaser and Yunpeng Men. OAJ acknowledges the support of Breakthrough Listen which is managed by the Breakthrough Prize Foundation. This work was performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government. JS acknowledges support from National Science Foundation grant AST-2205550.
Appendix A Detection of J0520-2553
As stated in section 4 there was a serendipitous detection of J0520-2553 (see fig. 5), using both the acceleration search and single burst search pipelines in S band. The fluxes were calculated using eq. 3 along with applying an offset from the boresight which is assumed to be Gaussian, $\exp\left(-\left({4\ln 2\cdot\theta^{2}}\right)/{\text{FWHM}^{2}}\right)$. Where $\theta$ is the angular separation from the boresight.
@@ Appendix A Detection of J0520-2553
Figure 5: Top: Pulse profile of J0520-2553 obtained from the accelsearch pipeline at a DM of 33.7 $\text{pc\leavevmode\nobreak\ cm}^{-3}$. Bottom: Single burst pulse also detected at 33.7 $\text{pc\leavevmode\nobreak\ cm}^{-3}$  in S-band.

### v2

@@ 5 Conclusions
acknowledgements
The authors acknowledge the insights and comments from Jules Halpern, Karen Perez, Wael Farah, Colin Clark, Scott Ransom, Lawrence Toomey, Joe Glaser and Yunpeng Men. OAJ acknowledges the support of the Trinty College Dublin, School of Physics and Breakthrough Listen which is managed by the Breakthrough Prize Foundation. This work was performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government. JS acknowledges support from National Science Foundation grant AST-2205550. We also thank the anonymous reviewer for their careful reading of our manuscript and their many insightful comments and suggestions.
6 appendix
6.1 Detection of J0520-2553
As stated in §4 there was a serendipitous detection of J0520-2553 (see Fig. 6), using both the acceleration search and single burst search pipelines in S band. The fluxes were calculated using eq. 3 along with applying an offset from the boresight which is assumed to be Gaussian, $\exp\left(-\left({4\ln 2\cdot\theta^{2}}\right)/{\text{FWHM}^{2}}\right)$. Where $\theta$ is the angular separation from the boresight.
Table 4: Flux limits in mJy for S-band observations of J0520-2553 attained from each of the search pipelines in this work.
@@ 6.1 Detection of J0520-2553
Figure 5: Flux limits in mJy for S-band observations of J0520-2553 from
Figure 6: Top: Pulse profile of J0520-2553 obtained from the accelsearch pipeline at a DM of 33.7 $\text{pc~cm}^{-3}$. Bottom: Single burst pulse also detected at 33.7 $\text{pc~cm}^{-3}$  in S-band.